carlos is 3 years older than juan. Jose is 4 years younger than juan. the sum of the three ages is 140. how old is juan?
Answer by addingup(3677)

Carlos: Juan + 3
Jose: Juan - 4
Let Juan be x. Then Carlos will be x + 3 and Jose x - 4
-------------------------------
x + x + 3 + x - 4 = 140
3x - 1 = 140
3x = 141
x = 47
----------------------------
So Juan is 47
Carlos: 47 + 3 = 50
Jose: 47-4 = 43
47 + 50 + 43 = 140 Correct